Simplify the given boolean expression

(~ this symbol represents NOT)

F = A + ~A*B + ~A*~B*C + ~A*~B*~C*D + ~A*~B*~C*~D*E
Answer #1

we have,

F = A + ~A*B + ~A*~B*C + ~A*~B*~C*D + ~A*~B*~C*~D*E

we can write,

\small F = A + \bar{A}B + \bar{A}\bar{B}C+\bar{A}\bar{B}\bar{C}D+\bar{A}\bar{B}\bar{C}\bar{D}E

\small F = A + \bar{A}B + \bar{A}\bar{B}C+\bar{A}\bar{B}\bar{C}\left (D+\bar{D}E \right )

Using ditributive for the last term law we can say that,

\small F = A + \bar{A}B + \bar{A}\bar{B}C+\bar{A}\bar{B}\bar{C}\left ((D+\bar{D})(D+E) \right )

we know that D + D' = 1 hence,

\small F = A + \bar{A}B + \bar{A}\bar{B}C+\bar{A}\bar{B}\bar{C}\left ((1)(D+E) \right )

\small F = A + \bar{A}B + \bar{A}\bar{B}C+\bar{A}\bar{B}\bar{C}\left (D+E \right )

\small F = A + \bar{A}B + \bar{A}\bar{B}\left (C+\bar{C} (D+E ) \right )

Using ditributive for the last term law we can say that,

\small F = A + \bar{A}B + \bar{A}\bar{B}((C+\bar{C})(C+D+E))

we know that C + C' = 1 hence,

\small F = A + \bar{A}B + \bar{A}\bar{B}((1)(C+D+E))

\small F = A + \bar{A}B + \bar{A}\bar{B}(C+D+E)

\small F = A + \bar{A}(B + \bar{B}(C+D+E))

Using ditributive for the last term law we can say that,

\small F = A + \bar{A}((B + \bar{B})(B+C+D+E))

we know that B + B' = 1 hence,

\small F = A + \bar{A}((1)(B+C+D+E))

\small F = A + \bar{A}(B+C+D+E)

Using ditributive law we can say that,

\small F = (A + \bar{A})(A+B+C+D+E)

we know that A + A' = 1 hence,

\small F =(1)(A+B+C+D+E)

\small F =(A+B+C+D+E)

Hence we can write,

\small F =A+B+C+D+E